On earth, tin has 10 different stable isotopes. The heaviest, \ce{^{124}Sn}

X


124

​X

2


2

124

​Sn makes up 5.80\%

5.80% of naturally occurring tin atoms. How many atoms of \ce{^{124}Sn}

X


124

​X

2


2

124

​Sn are present in 18.3

18.3 g of a natural sample of tin?

Considering \ce{^{124}Sn}

X


124

​X

2


2

124

​Sn has an isotopic mass of \ce{{123.905} amu}

123.905 amu , calculate the mass of this isotope in the sample.


1
Expert's answer
2021-09-20T06:01:27-0400

The average atomic mass of Sn is 118.71 g/mol


the percentage of heaviest Sn is 5.80%


the given mass of Sn is 82g


The total moles of Sn will be = mass / atomic mass = 82/118.71=0.691


Total atoms of Sn in 82g = 6.023× 1023× 0.691= 4.16×1023


the percentage of heaviest Sn is 5.80%


So the total atoms of Sn124 = 5.80% X 4.16× 1023


Total atoms of Sn124 = 2.41× 1022 atoms


the mass of Sn124 will be ="\frac{2.41× 10^{22} ×124}{6.023×10^{23}}" = 4.96g